# Monotonicity properties of weighted geometric symmetrizations

### v1

@@ 3. New results
Proof.
Assume $0\leq\alpha_{1}<\alpha_{2}\leq\frac{1}{2}$ and let $\alpha=\frac{\alpha_{1}+\alpha_{2}-1}{2\alpha_{1}-1}$. Then $\alpha\in(0,1)$ and for every positive kernel operator $K$ on $L$ we have $S_{\alpha_{2}}(K)=S_{\alpha}(S_{\alpha_{1}}(K))$. Indeed, the kernel of the operator $S_{\alpha}(S_{\alpha_{1}}(K))$is equal to
$(s_{\alpha_{1}}(k)(x,y))^{\alpha}(s_{\alpha_{1}}(k)(y,x))^{1-\alpha}$
$=(k(x,y)^{\alpha_{1}}k(y,x)^{1-\alpha_{1}})^{\alpha}(k(y,x)^{\alpha_{1}}k(x,y)^{1-\alpha_{1}})^{1-\alpha}$
$=k(x,y)^{\alpha_{1}\alpha+(1-\alpha_{1})(1-\alpha)}k(y,x)^{\alpha(1-\alpha_{1})+\alpha_{1}(1-\alpha)}=k(x,y)^{\alpha_{2}}k(y,x)^{1-\alpha_{2}},$
which is a kernel of the operator $S_{\alpha_{2}}(K)$ since
$\alpha_{1}\alpha+(1-\alpha_{1})(1-\alpha)=\alpha(2\alpha_{1}-1)+1-\alpha_{1}=\alpha_{1}+\alpha_{2}-1+1-\alpha_{1}=\alpha_{2}$
and
$\alpha(1-\alpha_{1})+\alpha_{1}(1-\alpha)=\alpha(1-2\alpha_{1})+\alpha_{1}=1-\alpha_{1}-\alpha_{2}+\alpha_{1}=1-\alpha_{2}.$
It follows from (3.1) that
$\rho_{n}(\alpha_{2})=\sqrt{\rho(S_{\alpha_{2}}(K_{1})S_{\alpha_{2}}(K_{2})\cdots S_{\alpha_{2}}(K_{n}))\rho(S_{\alpha_{2}}(K_{n})S_{\alpha_{2}}(K_{n-1})\cdots S_{\alpha_{2}}(K_{1}))}$
$=\sqrt{\rho(S_{\alpha}(S_{\alpha_{1}}(K_{1}))\cdots S_{\alpha}(S_{\alpha_{1}}(K_{n})))\rho(S_{\alpha}(S_{\alpha_{1}}(K_{n}))\cdots S_{\alpha}(S_{\alpha_{1}}(K_{1})))}\leq$
$\sqrt{\rho(S_{\alpha_{1}}(K_{1})\cdots S_{\alpha_{1}}(K_{n}))^{\alpha}\rho(S_{\alpha_{1}}(K_{n})\cdots S_{\alpha_{1}}(K_{1}))^{1-\alpha}}\times$
$\sqrt{\rho(S_{\alpha_{1}}(K_{n})\cdots S_{\alpha_{1}}(K_{1}))^{\alpha}\rho(S_{\alpha_{1}}(K_{1})\cdots S_{\alpha_{1}}(K_{n}))^{1-\alpha}}$
$=\sqrt{\rho(S_{\alpha_{1}}(K_{1})\cdots S_{\alpha_{1}}(K_{n}))\rho(S_{\alpha_{1}}(K_{n})\cdots S_{\alpha_{1}}(K_{1}))}=\rho_{n}(\alpha_{1}),$
which proves that $\rho_{n}$ is decreasing on $[0,\frac{1}{2}]$.
Similarly, in the case $\frac{1}{2}\leq\alpha_{1}<\alpha_{2}\leq 1$ let $\alpha=\frac{\alpha_{1}+\alpha_{2}-1}{2\alpha_{2}-1}$. It follows that $\alpha\in(0,1)$ and $S_{\alpha_{1}}(K)=S_{\alpha}(S_{\alpha_{2}}(K))$ for every positive kernel operator on $L$. Similarly as before this holds since the kernel of $S_{\alpha}(S_{\alpha_{2}}(K))$ equals
$(s_{\alpha_{2}}(k)(x,y))^{\alpha}(s_{\alpha_{2}}(k)(y,x))^{1-\alpha}$
$=(k(x,y)^{\alpha_{2}}k(y,x)^{1-\alpha_{2}})^{\alpha}(k(y,x)^{\alpha_{2}}k(x,y)^{1-\alpha_{2}})^{1-\alpha}$
$=k(x,y)^{\alpha_{2}\alpha+(1-\alpha_{2})(1-\alpha)}k(y,x)^{\alpha(1-\alpha_{2})+\alpha_{2}(1-\alpha)}=k(x,y)^{\alpha_{1}}k(y,x)^{1-\alpha_{1}},$
which is a kernel of the operator $S_{\alpha_{1}}(K)$ since $\alpha_{2}\alpha+(1-\alpha_{2})(1-\alpha)=\alpha_{1}$ and $\alpha(1-\alpha_{2})+\alpha_{2}(1-\alpha)=1-\alpha_{1}.$ From (3.1) we obtain
$\rho_{n}(\alpha_{1})=\sqrt{\rho(S_{\alpha_{1}}(K_{1})\cdots S_{\alpha_{1}}(K_{n}))\rho(S_{\alpha_{1}}(K_{n})\cdots S_{\alpha_{1}}(K_{1}))}$
$=\sqrt{\rho(S_{\alpha}(S_{\alpha_{2}}(K_{1}))\cdots S_{\alpha}(S_{\alpha_{2}}(K_{n})))\rho(S_{\alpha}(S_{\alpha_{2}}(K_{n}))\cdots S_{\alpha}(S_{\alpha_{2}}(K_{1})))}\leq$
$\sqrt{\rho(S_{\alpha_{2}}(K_{1})\cdots S_{\alpha_{2}}(K_{n}))^{\alpha}\rho(S_{\alpha_{2}}(K_{n})\cdots S_{\alpha_{2}}(K_{1}))^{1-\alpha}}\times$
$\sqrt{\rho(S_{\alpha_{2}}(K_{n})\cdots S_{\alpha_{2}}(K_{1}))^{\alpha}\rho(S_{\alpha_{2}}(K_{1})\cdots S_{\alpha_{2}}(K_{n}))^{1-\alpha}}$
$=\sqrt{\rho(S_{\alpha_{2}}(K_{1})\cdots S_{\alpha_{2}}(K_{n}))\rho(S_{\alpha_{2}}(K_{n})\cdots S_{\alpha_{2}}(K_{1}))}=\rho_{n}(\alpha_{2}),$
which proves that $\rho_{n}$ is increasing on $[\frac{1}{2},1]$. ∎
In the case $n=1$ we obtain the result from [6, Theorem 2.7].
Corollary 3.2.
Let $K$ be a positive kernel operator on $L^{2}(X,\mu)$ and
$\rho\in\{r,r_{ess},\gamma,\|\cdot\|,w\}$. Then a function $\rho_{1}:[0,1]\to[0,\infty)$, defined by $\rho_{1}(\alpha)=\rho(S_{\alpha}(K))$, is decreasing on $[0,\frac{1}{2}]$ and increasing on $[\frac{1}{2},1]$.
In particular, $\rho(S_{\alpha}(K))\geq\rho(S(K))$ for each $\alpha\in[0,1]$.
Corollary 3.3.
Let $K_{1}$ and $K_{2}$ be positive kernel operators on $L^{2}(X,\mu)$ and $\rho\in\{r,r_{ess}\}$. Then $\rho_{2}(\alpha)=\rho(S_{\alpha}(K_{1})S_{\alpha}(K_{2}))$ is decreasing on $[0,\frac{1}{2}]$ and increasing on $[\frac{1}{2},1]$.
In particular, $\rho(S_{\alpha}(K_{1})S_{\alpha}(K_{2}))\geq\rho(S(K_{1})S(K_{2}))$ for each $\alpha\in[0,1]$.
Proof.
The statement follows from Theorem 3.1 since $\rho(AB)=\rho(BA)$ for all bounded operators $A$ and $B$ on $L^{2}(X,\mu)$. ∎
Proposition 3.4.
Let $K_{1},\ldots,K_{n}$ be positive kernel operators on $L=L^{2}(X,\mu)$ and $\rho\in\{r,r_{ess},\gamma,\|\cdot\|,w\}$. Then a function $\widetilde{\rho_{n}}:[0,1]\to[0,\infty)$, defined by $\widetilde{\rho_{n}}(\alpha)=\rho(S_{\alpha}(K_{1})+\cdots+S_{\alpha}(K_{n}))$, is decreasing on $[0,\frac{1}{2}]$ and increasing on $[\frac{1}{2},1]$.
Proof.
Let $0\leq\alpha_{1}<\alpha_{2}\leq\frac{1}{2}$. For $\alpha=\frac{\alpha_{1}+\alpha_{2}-1}{2\alpha_{1}-1}$ we have $\alpha\in(0,1)$ and $S_{\alpha_{2}}(K)=S_{\alpha}(S_{\alpha_{1}}(K))$ (see the proof of Theorem 3.1). By applying (3.2) we obtain
$\widetilde{\rho_{n}}(\alpha_{2})=\rho(S_{\alpha_{2}}(K_{1})+\cdots+S_{\alpha_{2}}(K_{n}))=\rho(S_{\alpha}(S_{\alpha_{1}}(K_{1}))+\ldots+S_{\alpha}(S_{\alpha_{1}}(K_{n})))$
$\leq\rho(S_{\alpha_{1}}(K_{1})+\cdots+S_{\alpha_{1}}(K_{n}))=\widetilde{\rho_{n}}(\alpha_{1}),$
which proves that $\widetilde{\rho_{n}}$ is decreasing on $[0,\frac{1}{2}]$.
For $\frac{1}{2}\leq\alpha_{1}<\alpha_{2}\leq 1$ let $\alpha=\frac{\alpha_{1}+\alpha_{2}-1}{2\alpha_{2}-1}$. Then $\alpha\in(0,1)$ and $S_{\alpha_{1}}(K)=S_{\alpha}(S_{\alpha_{2}}(K))$ (see the proof of Theorem 3.1). By (3.2) it follows that
$\widetilde{\rho_{n}}(\alpha_{1})=\rho(S_{\alpha_{1}}(K_{1})+\cdots+S_{\alpha_{1}}(K_{n}))=\rho(S_{\alpha}(S_{\alpha_{2}}(K_{1}))+\ldots+S_{\alpha}(S_{\alpha_{2}}(K_{n})))$
$\leq\rho(S_{\alpha_{2}}(K_{1})+\cdots+S_{\alpha_{2}}(K_{n}))=\widetilde{\rho_{n}}(\alpha_{2}),$
which proves that $\widetilde{\rho_{n}}$ is increasing on $[\frac{1}{2},1]$. ∎
By applying (2.13) also the following more general result follows.
@@
Let $\alpha\in[0,1]$. By (2.13) it follows that
$S_{\alpha}(K_{i1})+\cdots+S_{\alpha}(K_{im})\leq S_{\alpha}(K_{i1}+\cdots K_{im})$ (3.4)
holds for all $i=1,\ldots,n$. From (3.4) and (3.1) it follows that
$\overline{\rho}_{n}(\alpha)\leq\rho\left(S_{\alpha}(K_{11}+\cdots+K_{1m})\cdots S_{\alpha}(K_{n1}+\cdots+K_{nm})\right)^{\frac{1}{2}}\times$
$\rho\left(S_{\alpha}(K_{n1}+\cdots+K_{nm})\cdots S_{\alpha}(K_{11}+\cdots+K_{1m})\right)^{\frac{1}{2}}$
$\leq\rho(\left(K_{11}+\cdots+K_{1m})\cdots(K_{n1}+\cdots+K_{nm})\right)^{\frac{\alpha}{2}}\times$
$\rho(\left(K_{n1}+\cdots+K_{nm})\cdots(K_{11}+\cdots+K_{1m})\right)^{\frac{1-\alpha}{2}}\times$
$\rho(\left(K_{n1}+\cdots+K_{nm})\cdots(K_{11}+\cdots+K_{1m})\right)^{\frac{\alpha}{2}}\times$
$\rho(\left(K_{11}+\cdots+K_{1m})\cdots(K_{n1}+\cdots+K_{nm})\right)^{\frac{1-\alpha}{2}}$
$=\rho(\left(K_{11}+\cdots+K_{1m})\cdots(K_{n1}+\cdots+K_{nm})\right)^{\frac{1}{2}}\times$
$\rho(\left(K_{n1}+\cdots+K_{nm})\cdots(K_{11}+\cdots+K_{1m})\right)^{\frac{1}{2}},$
which proves (3.3).
To prove that $\overline{\rho}_{n}$ is decreasing on $[0,\frac{1}{2}]$ let $0\leq\alpha_{1}<\alpha_{2}\leq\frac{1}{2}$. For $\alpha=\frac{\alpha_{1}+\alpha_{2}-1}{2\alpha_{1}-1}$ we have $\alpha\in(0,1)$ and $S_{\alpha_{2}}(K)=S_{\alpha}(S_{\alpha_{1}}(K))$. Then by (3.3)
$\overline{\rho}_{n}(\alpha_{2})=$
$\left(\rho(\left(S_{\alpha}(S_{\alpha_{1}}(K_{11}))+\cdots+S_{\alpha}(S_{\alpha_{1}}(K_{1m}))\right)\cdots(S_{\alpha}(S_{\alpha_{1}}(K_{n1}))+\cdots+S_{\alpha}((S_{\alpha_{1}}(K_{nm})))\right)^{\frac{1}{2}}\times$
$\left(\rho(\left(S_{\alpha}(S_{\alpha_{1}}(K_{n1}))+\cdots+S_{\alpha}(S_{\alpha_{1}}(K_{nm}))\right)\cdots(S_{\alpha}(S_{\alpha_{1}}(K_{11}))+\cdots+S_{\alpha}((S_{\alpha_{1}}(K_{1m})))\right)^{\frac{1}{2}}$
$\leq\left(\rho((S_{\alpha_{1}}(K_{11})+\cdots+S_{\alpha_{1}}(K_{1m}))\cdots(S_{\alpha_{1}}(K_{n1})+\cdots+S_{\alpha_{1}}(K_{nm})))\right)^{\frac{1}{4}}\times$
$\left(\rho((S_{\alpha_{1}}(K_{n1})+\cdots+S_{\alpha_{1}}(K_{nm}))\cdots(S_{\alpha_{1}}(K_{11})+\cdots+S_{\alpha_{1}}(K_{1m})))\right)^{\frac{1}{4}}\times$
$\left(\rho((S_{\alpha_{1}}(K_{n1})+\cdots+S_{\alpha_{1}}(K_{nm}))\cdots(S_{\alpha_{1}}(K_{11})+\cdots+S_{\alpha_{1}}(K_{1m})))\right)^{\frac{1}{4}}\times$
$\left(\rho((S_{\alpha_{1}}(K_{11})+\cdots+S_{\alpha_{1}}(K_{1m}))\cdots(S_{\alpha_{1}}(K_{n1})+\cdots+S_{\alpha_{1}}(K_{nm})))\right)^{\frac{1}{4}}=\overline{\rho}_{n}(\alpha_{1}),$
which establishes that $\overline{\rho}_{n}$ is decreasing on $[0,\frac{1}{2}]$.
To prove that $\overline{\rho}_{n}$ is increasing on $[\frac{1}{2},1]$ let $\frac{1}{2}\leq\alpha_{1}<\alpha_{2}\leq 1$. For $\alpha=\frac{\alpha_{1}+\alpha_{2}-1}{2\alpha_{2}-1}$ we have $\alpha\in(0,1)$ and $S_{\alpha_{1}}(K)=S_{\alpha}(S_{\alpha_{2}}(K))$. Similarly as above it follows from (3.3) that $\overline{\rho}_{n}(\alpha_{1})\leq\overline{\rho}_{n}(\alpha_{2})$, which completes the proof.

### v2

@@ 2. Preliminaries
In the next section we generalize the above two results to the setting of bounded sets of positive kernel operators on $L^{2}(X,\mu)$ (Theorems 3.4 and 3.10 below).
3. New results
In this section we extend Theorems 2.4 and 2.5 to the setting of bounded sets of positive kernel operators on $L^{2}(X,\mu)$.
Let $\Psi$ be a bounded set of positive kernel operators on $L=L^{2}(X,\mu)$ and $\alpha\in[0,1]$. Denote by $\Psi^{*}$ and $S_{\alpha}(\Psi)$ bounded sets of positive kernel operators on $L$ defined by $\Psi^{*}=\{A^{*}:A\in\Psi\}$ and $S_{\alpha}(\Psi)=\Psi^{(\alpha)}\circ(\Psi^{*})^{(1-\alpha)}=\{A^{(\alpha)}\circ(B^{*})^{(1-\alpha)}:A,B\in\Psi\}$. Note that $S_{\alpha}(\Psi)^{*}=S_{\alpha}(\Psi^{*})=\{(A^{*})^{(\alpha)}\circ B^{(1-\alpha)}:A,B\in\Psi\}$.
Let $\Psi,\Psi_{1},\ldots,\Psi_{n}$ be bounded sets of positive kernel operators on $L=L^{2}(X,\mu)$, $\rho\in\{r,\hat{r},r_{ess},\hat{r}_{ess}\}$ and $\alpha\in[0,1]$. In [7, Proposition 4.2] the following generalizations of (2.24) and (2.25) were proved
$\rho(S_{\alpha}(\Psi_{1})\cdots S_{\alpha}(\Psi_{n}))\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;$
$\leq\rho\left((\Psi_{1}\cdots\Psi_{n})^{(\alpha)}\circ((\Psi_{n}\cdots\Psi_{1})^{*})^{(1-\alpha)}\right)\leq\rho(\Psi_{1}\cdots\Psi_{n})^{\alpha}\,\rho(\Psi_{n}\cdots\Psi_{1})^{1-\alpha}$ (3.1)
and
$\rho(S_{\alpha}(\Psi_{1})+\cdots+S_{\alpha}(\Psi_{m}))\leq\rho\left(S_{\alpha}(\Psi_{1}+\cdots+\Psi_{m})\right)$ (3.2)
$\leq$ $\rho(\Psi_{1}+\cdots+\Psi_{m}).$
In particular ([7, Proposition 4.2 (58)]),
$\rho(S_{\alpha}(\Psi))\leq\rho(\Psi).$ (3.3)
Let us denote $\|\Psi\|=\sup_{A\in\Psi}\|A\|$, $\gamma(\Psi)=\sup_{A\in\Psi}\gamma(A)$ and $w(\Psi)=\sup_{A\in\Psi}w(A)$.
We show below in Corollary 3.3 that (3.1) and (3.2) hold also for all $\rho\in\{\|\cdot\|,\gamma,w\}$. To do this we first point out the following analogue of Theorem 2.3 for all $\rho\in\{\|\cdot\|,\gamma,w\}$ that easily follows from Theorem 2.1 and (2.21).
Corollary 3.1.
Let $\{\Psi_{ij}\}_{i=1,j=1}^{k,m}$ be bounded sets of positive kernel operators on a Banach function space $L$ and assume that $\alpha_{1},\ldots,\alpha_{m}$ are positive numbers such that $\sum_{i=1}^{m}\alpha_{i}=1$. Then (2.22) and (2.23) hold for $\rho=\|\cdot\|$.
If, in addition, $L$ and $L^{*}$ have order continuous norms, then (2.22) and (2.23) hold also for $\rho=\gamma$.
If, in addition, $L=L^{2}(X,\mu)$, then (2.22) and (2.23) hold also for $\rho=w$.
@@ 3. New results
Inequalities (2.22) in all three cases follow from Theorem 2.1, while inequalities (2.23) follow from (2.21), monotonicity of $\rho$ and from Theorem 2.2. ∎
Remark 3.2.
Let $\mathcal{L}$ denote the set of all Banach sequence spaces $L$ such that all standard vectors are included in $L$ and have norm $1$ (for precise definitions see e.g. [7], [22], [32]). If $\alpha_{1},\ldots,\alpha_{m}$ are positive numbers such that $\sum_{i=1}^{m}\alpha_{i}\geq 1$, then for $\rho=\|\cdot\|$ inequalities (2.22) and (2.23) hold also for bounded sets of nonnegative matrices that define positive operators on $L\in\mathcal{L}$ by [22, Theorem 2.1 (ii)]. If, in addition, $L$ and $L^{*}$ have order continuous norms then the same conclusion holds also for $\rho=\gamma$ by [22, Theorem 3.5]. However, if in addition, $L=l^{2}$, then the same conclusion does not hold for $\rho=w$ (see e.g. [13, Example 3.2]).
The following result establishes an analogue of (3.1) and (3.2) for all $\rho\in\{\|\cdot\|,\gamma,w\}$ and it extends [6, Proposition 2.2].
@@
Let $\Psi_{1},\ldots,\Psi_{m}$ be bounded sets of positive kernel operators on $L^{2}(X,\mu)$ and let $\alpha\in[0,1]$. Then (3.1) and (3.2) hold for all $\rho\in\{\|\cdot\|,\gamma,w\}$.
Proof.
Let $\rho\in\{\|\cdot\|,\gamma,w\}$. By Corollary 3.1 we have
$\rho\left(S_{\alpha}(\Psi_{1})\cdots S_{\alpha}(\Psi_{m})\right)=\rho\left((\Psi_{1}^{(\alpha)}\circ(\Psi_{1}^{*})^{(1-\alpha)})\cdots\left(\Psi_{m}^{(\alpha)}\circ(\Psi_{m}^{*})^{(1-\alpha)}\right)\right)$
$\leq$ $\rho\left((\Psi_{1}\cdots\Psi_{m})^{(\alpha)}\circ((\Psi_{m}\cdots\Psi_{1})^{*})^{(1-\alpha)}\right)$
$\leq$ $\rho(\Psi_{1}\cdots\Psi_{m})^{\alpha}\,\rho((\Psi_{m}\cdots\Psi_{1})^{*})^{1-\alpha}=\rho(\Psi_{1}\cdots\Psi_{m})^{\alpha}\,\rho(\Psi_{m}\cdots\Psi_{1})^{1-\alpha},$
where the last equality follows from the fact that $\rho(\Psi)=\rho(\Psi^{*})$. This completes the proof of (3.1) for $\rho\in\{\|\cdot\|,\gamma,w\}$. The inequalities in (3.2) are proved in similar way by applying (2.23). ∎
The following results generalizes Theorem 2.4.
Theorem 3.4.
Let $\Psi_{1},\ldots,\Psi_{n}$ be bounded sets of positive kernel operators on $L=L^{2}(X,\mu)$ and $\rho\in\{r,\hat{r},r_{ess},\hat{r}_{ess},\|\cdot\|,\gamma,w\}$. Define $\rho_{n}:[0,1]\to[0,\infty)$ by
$\rho_{n}(\alpha)=\sqrt{\rho(S_{\alpha}(\Psi_{1})S_{\alpha}(\Psi_{2})\cdots S_{\alpha}(\Psi_{n}))\rho(S_{\alpha}(\Psi_{n})S_{\alpha}(\Psi_{n-1})\cdots S_{\alpha}(\Psi_{1}))}.$
Proof.
Let $0\leq\alpha_{1}<\alpha_{2}\leq\frac{1}{2}$ and $\alpha=\frac{\alpha_{1}+\alpha_{2}-1}{2\alpha_{1}-1}$. Then $\alpha\in(0,1)$ and
$S_{\alpha_{2}}(\Psi)\subset S_{\alpha}(S_{\alpha_{1}}(\Psi))$ (3.4)
for every bounded set of positive kernel operators on $L=L^{2}(X,\mu)$. Indeed, for $A^{(\alpha_{2})}\circ(B^{*})^{(1-\alpha_{2})}\in S_{\alpha_{2}}(\Psi)$ where $A,B\in\Psi$ we have
$A^{(\alpha_{2})}\circ(B^{*})^{(1-\alpha_{2})}=A^{(\alpha\alpha_{1}+(1-\alpha_{1})(1-\alpha))}\circ(B^{*})^{(\alpha(1-\alpha_{1})+\alpha_{1}(1-\alpha))}$
$=(A^{(\alpha_{1})}\circ(B^{*})^{(1-\alpha_{1})})^{(\alpha)}\circ((B^{*})^{(\alpha_{1})}\circ A^{(1-\alpha_{1})})^{(1-\alpha)}\in S_{\alpha}(S_{\alpha_{1}}(\Psi)),$
since $A^{(\alpha_{1})}\circ(B^{*})^{(1-\alpha_{1})}\in S_{\alpha_{1}}(\Psi)$ and $(B^{*})^{(\alpha_{1})}\circ A^{(1-\alpha_{1})}\in S_{\alpha_{1}}(\Psi^{*})=S_{\alpha_{1}}(\Psi)^{*}$. From (3.4) and (3.1) it follows that
$\rho_{n}(\alpha_{2})=\sqrt{\rho(S_{\alpha_{2}}(\Psi_{1})\cdots S_{\alpha_{2}}(\Psi_{n}))\rho(S_{\alpha_{2}}(\Psi_{n})\cdots S_{\alpha_{2}}(\Psi_{1}))}$
$\leq\sqrt{\rho(S_{\alpha}(S_{\alpha_{1}}(\Psi_{1}))\cdots S_{\alpha}(S_{\alpha_{1}}(\Psi_{n})))\rho(S_{\alpha}(S_{\alpha_{1}}(\Psi_{n}))\cdots S_{\alpha}(S_{\alpha_{1}}(\Psi_{1})))}$
$\leq\sqrt{\rho{((S_{\alpha_{1}}(\Psi_{1})\cdots S_{\alpha_{1}}(\Psi_{n}))^{\alpha}}\rho{((S_{\alpha_{1}}(\Psi_{n})\cdots S_{\alpha_{1}}(\Psi_{1}))^{1-\alpha}}}\times$
$\sqrt{\rho{((S_{\alpha_{1}}(\Psi_{n})\cdots S_{\alpha_{1}}(\Psi_{1}))^{\alpha}}\rho{((S_{\alpha_{1}}(\Psi_{1})\cdots S_{\alpha_{1}}(\Psi_{n}))^{1-\alpha}}}$
$=\sqrt{\rho(S_{\alpha_{1}}(\Psi_{1})\cdots S_{\alpha_{1}}(\Psi_{n}))\rho(S_{\alpha_{1}}(\Psi_{n})\cdots S_{\alpha_{1}}(\Psi_{1}))}=\rho_{n}(\alpha_{1}),$
which proves that $\rho_{n}$ is a decreasing function on $[0,\frac{1}{2}]$.
To prove the case $\frac{1}{2}\leq\alpha_{1}<\alpha_{2}\leq 1$, let $\alpha=\frac{\alpha_{1}+\alpha_{2}-1}{2\alpha_{2}-1}$. Then $\alpha\in(0,1)$ and
$S_{\alpha_{1}}(\Psi)\subset S_{\alpha}(S_{\alpha_{2}}(\Psi))$ (3.5)
for every bounded set $\Psi$ of positive kernel operators on $L=L^{2}(X,\mu)$ Indeed, let $C^{(\alpha_{1})}\circ(D^{*})^{(1-\alpha_{1})}\in S_{\alpha_{1}}(\Psi)$, where $C,D\in\Psi$. Then
$C^{(\alpha_{1})}\circ(D^{*})^{(1-\alpha_{1})}=C^{(\alpha\alpha_{2}+(1-\alpha_{2})(1-\alpha))}\circ(D^{*})^{(\alpha(1-\alpha_{2})+\alpha_{2}(1-\alpha))}$
$=(C^{(\alpha_{2})}\circ(D^{*})^{(1-\alpha_{2})})^{(\alpha)}\circ((D^{*})^{(\alpha_{2})}\circ C^{(1-\alpha_{2})})^{(1-\alpha)}\in S_{\alpha}(S_{\alpha_{2}}(\Psi)),$
since $C^{(\alpha_{2})}\circ(D^{*})^{(1-\alpha_{2})}\in S_{\alpha_{2}}(\Psi)$ and $(D^{*})^{(\alpha_{2})}\circ C^{(1-\alpha_{2})}\in S_{\alpha_{2}}(\Psi^{*})=S_{\alpha_{2}}(\Psi)^{*}$. Applying (3.5) and (3.1) we obtain
$\rho_{n}(\alpha_{1})=\sqrt{\rho(S_{\alpha_{1}}(\Psi_{1})\cdots S_{\alpha_{1}}(\Psi_{n}))\rho(S_{\alpha_{1}}(\Psi_{n})\cdots S_{\alpha_{1}}(\Psi_{1}))}$
$\leq\sqrt{\rho(S_{\alpha}(S_{\alpha_{2}}(\Psi_{1}))\cdots S_{\alpha}(S_{\alpha_{2}}(\Psi_{n})))\rho(S_{\alpha}(S_{\alpha_{2}}(\Psi_{n}))\cdots S_{\alpha}(S_{\alpha_{2}}(\Psi_{1})))}$
$\leq\sqrt{\rho{((S_{\alpha_{2}}(\Psi_{1})\cdots S_{\alpha_{2}}(\Psi_{n}))^{\alpha}}\rho{((S_{\alpha_{2}}(\Psi_{n})\cdots S_{\alpha_{2}}(\Psi_{1}))^{1-\alpha}}}\times$
$\sqrt{\rho{((S_{\alpha_{2}}(\Psi_{n})\cdots S_{\alpha_{2}}(\Psi_{1}))^{\alpha}}\rho{((S_{\alpha_{2}}(\Psi_{1})\cdots S_{\alpha_{2}}(\Psi_{n}))^{1-\alpha}}}$
$=\sqrt{\rho(S_{\alpha_{2}}(\Psi_{1})\cdots S_{\alpha_{2}}(\Psi_{n}))\rho(S_{\alpha_{2}}(\Psi_{n})\cdots S_{\alpha_{2}}(\Psi_{1}))}=\rho_{n}(\alpha_{2}),$
which completes the proof. ∎
The following corollary generalizes [6, Theorem 2.7] ([7, Corollary 3.2]).
Corollary 3.5.
Let $\Psi$ be a bounded set of positive kernel operators on $L=L^{2}(X,\mu)$ and $\rho\in\{r,\hat{r},r_{ess},\hat{r}_{ess},\|\cdot\|,\gamma,w\}$. Then the function $\rho_{1}:[0,1]\to[0,\infty)$, defined by $\rho_{1}(\alpha)=\rho(S_{\alpha}(\Psi))$, is decreasing on $[0,\frac{1}{2}]$ and increasing on $[\frac{1}{2},1]$.
In particular, $\rho(S_{\alpha}(\Psi))\geq\rho(S(\Psi))$ for each $\alpha\in[0,1]$.
The following corollary generalizes [8, Corollary 3.3].
Corollary 3.6.